6y + 8x = -432
i dont know how to find the x int. and the y int. or the slope in this problem

Answer by tutorcecilia(2144)

The values for the x-intercept = (x,0)
The values for the y-intercept = (0,b)
6y + 8x = -432
.
To find the x-intercept, let y=0 and solve for x
6y+8x=-432
6(0)+8x=-423
8x=-432
x=-54
x-intercept = (-54, 0)
.
To find the y-intercept, let x=0 and solve for y
6y+8(0)=-432
6y=-432
y=-72
y-intercept = (0, -72)
.